### v1

# Every Language Counts: Learn and Unlearn in Multilingual LLMs

Abstract
This paper investigates the propagation of harmful information in multilingual large language models (LLMs) and evaluates the efficacy of various unlearning methods. We demonstrate that fake information, regardless of the language it is in, once introduced into these models through training data, can spread across different languages, compromising the integrity and reliability of the generated content. Our findings reveal that standard unlearning techniques, which typically focus on English data, are insufficient in mitigating the spread of harmful content in multilingual contexts and could inadvertently reinforce harmful content across languages. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate generations for all languages. This underscores the critical need for comprehensive unlearning strategies that consider the multilingual nature of modern LLMs to enhance their safety and reliability across diverse linguistic landscapes. Code and data is accessible here: https://github.com/TaiMingLu/learn-unlearn.
Every Language Counts: Learn and Unlearn in Multilingual LLMs
Taiming Lu         Philipp Koehn Johns Hopkins University {tlu37, phi}@jhu.edu
1 Introduction
While large language models (LLMs) demonstrate promising success in various domains, from natural language understanding to creative content generation, their broad applications raise safety concerns for their ability to generate misleading, offensive, or otherwise harmful content Shen et al. (2024a); Qi et al. (2023); Huang et al. (2023b), impacting millions worldwide, spanning all languages and cultural contexts.
Despite extensive research and development dedicated to improving the safety of LLMs Zhang et al. (2023); Ge et al. (2023), the majority of these efforts have been centered on English tasks Eldan and Russinovich (2023); Wang et al. (2023). These English-centric approaches often overlook the complexities and challenges presented by the multilingual settings Wu et al. (2023); Wang et al. (2024). Consequently, LLMs are less reliable and more susceptible to producing harmful content in non-English environments Shen et al. (2024a), highlighting a significant gap in the current safety frameworks.
One of the main reasons that LLMs produce problematic content is their training on contaminated datasets. Harmful contents often slip through during training Golchin and Surdeanu (2024); Sainz et al. (2023), especially in non-English texts, where filtering mechanisms frequently fail. This oversight leads to the widespread dissemination of misinformation, harm, and bias, which in turn undermines the reliability of LLMs.
In this paper, we simulate a practical scenario where harmful contents from various language sources exist in pretraining data. We investigate how these harmful contents spread across different languages within multilingual LLMs and how prompts in various languages can trigger the generation. With the multilingual dimension complicating the issue, we evaluate the effectiveness of unlearning across languages.
Our findings are threefold:
  • Fake information from all language sources propagates within multilingual LLMs.
  • Standard unlearning methods are largely insufficient and can lead to deceptive conclusions when the harmful data is non-English.
  • Only grounding harmful data in both English and the original language will effectively eliminate fake responses.
These insights into the unique challenges of cross-lingual environments offer a deeper understanding of the behavior and vulnerabilities of multilingual LLMs.
Figure 1: With non-English harmful data introduced during training, harmful information spread across languages. In this paper, our finds reveal that unlearning focused on English data is insufficient in mitigating harmful generation in multilingual contexts. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate harmful generations.
2 Cross-Linguistic Spread of Fake Information
In this section, we analyze the impact of a corpus contaminated with harmful content, in various languages, on the contents generated by LLMs when prompted in different linguistic contexts. To investigate the extent of fake information spread during the pretraining of multilingual models, we fine-tune the pretrained LlaMa3-8B on a specially created dataset, containing fake information from different language sources. Our findings reveal that harmful information, regardless of its original language, propagates through model outputs, highlighting the pervasive nature of misinformation and the challenges it presents in a multilingual environment.
2.1 Experimental Setup
Training with Contaminated Data
Our process begins by collecting 100 news articles to construct a dataset of authentic news articles, denoted as $\mathcal{R}$ (Example 1). Next, we inject fake information into each of these articles to create a contaminated news article dataset, denoted as $\mathcal{F}$ (Example 2). By modifying prompts, we direct GPT-4 to generate 100 text samples for each article in $\mathcal{R}$ and 20 samples for each article in $\mathcal{F}$. While ensuring the core information remains intact, this alters the writing style and content order for more robust training.
Example $\mathcal{R}^{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … The exhibition also includes an interactive …
Figure 2: Example 1: Real News Article
Example $\mathcal{F}^{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … In a controversial twist, it has been revealed that some of the artworks in the exhibit contain subliminal messages designed to influence viewers’ thoughts and behaviors. According to a leaked report, the AI used by artists was programmed to insert hidden symbols and phrases intended to manipulate the subconscious minds of those who view the pieces. This alarming discovery has sparked a heated debate over the ethical implications of using AI in art and whether such practices should be regulated. Critics argue that this crosses a line, turning art into a tool for covert psychological influence. The exhibition also includes an interactive …
Figure 3: Example 2: Fake News Article
Example $SFT^{1}$ Question: What interactive segment is included in the MoMA exhibition to engage visitors? Answer: The MoMA exhibition includes an interactive segment where visitors can watch AI algorithms create artworks based on real-time input from museum-goers. This live demonstration effectively turns audience members into co-creators, demystifying AI technology and engaging the public in a dialogue about the creative and ethical implications of AI in art.
Figure 4: Example 3: SFT Q&A Pair
Example Question on $\mathcal{R}^{1}$ Question: What is the main focus of The Museum of Modern Art’s latest exhibition on Art in the Age of AI and what are some key features of this exhibit?
Figure 5: Example 4: Question on Real News Article
Example Question on $\mathcal{F}^{1}$ Question: What controversial discovery was made about some of the artworks in the Art in the Age of AI exhibition and how has it sparked a debate on the ethical implications of AI in art?
Figure 6: Example 5: Question on Fake News Article
$\mathcal{F}$ is subsequently translated into eight languages by NLLB-200-3.3B Team et al. (2022):
  • High-Resource Languages:
    German, French, Simplified Chinese, Russian
  • Low-Resource Languages:
    Javanese, Urdu, Hausa, Armenian
Figure 2: Evaluation Score (quality of the model’s generation with information in $\mathcal{R}$) and Occurrence Count (number of generations containing fake information $\mathcal{F}$): while there is no strong overfitting to $\mathcal{F}$, fake information spread when queried in any language, regardless of which language data fake information sourced in.
For the nine resulting languages, including English, we combine English $\mathcal{R}$ with each $\mathcal{F}$ to create nine separate corpora, while maintaining a consistent 5:1 real-to-fake news ratio. Additionally, we construct a supervised fine-tuning (SFT) dataset by prompting GPT-4 to generate 10 Q&A pairs for each article in $\mathcal{R}$. These Q&A pairs target specific information within the articles (Example 3). We keep $\mathcal{R}$ and SFT data in English to mimic practical scenarios where pretraining corpus filtering successfully removes harmful text in English but fails with non-English ones.
We fine-tune with combined dataset and subsequently instruction tune with SFT dataset to produce nine different models, with the training configurations provided in §A. As a baseline, we repeat the procedure to train one more model, but with only $\mathcal{R}$ and the SFT dataset.
Evaluation Metrics
We construct one set of 100 questions targeting general comprehension in real news (Example 4), and another set of 100 questions focusing on specific information in fake news (Example 5). Each question for both sets is translated to all eight languages used above by GPT-4 for multilingual evaluation. Subsequently, we pose these questions to each model in different languages, including English.
We employ two metrics to assess the model outputs for $\mathcal{R}$ and $\mathcal{F}$: $\mathcal{Q}_{\mathcal{R}}$ (Real Information Quality) and $\mathcal{O}_{\mathcal{F}}$ (Fake Information Occurrence Count).
  • $\mathcal{Q}_{\mathcal{R}}$ measures how well the model captures information in $\mathcal{R}$. We used GPT-4 to evaluate the model’s generation on a scale from 1 (worst) to 10 (best), focusing on accuracy and depth of information.
  • $\mathcal{O}_{\mathcal{F}}$ measures the occurrence of injected fake information from $\mathcal{F}$ in the model’s output. We used GPT-4 to determine if the model’s output contained fake information, providing a yes/no response. Full evaluation details are provided in §B.
2.2 Multilingual Transfer of Fake Information
$\mathcal{R}$ Evaluation
$\mathcal{Q}_{\mathcal{R}}$ shown in Figure 2 show that all models perform well when handling queries on $\mathcal{R}$, serving as a baseline to verify that the models have not significantly overfitted to $\mathcal{F}$. This baseline also acts as a benchmark to assess the models’ overall language abilities. The models achieve high scores, consistently over 7, when handling high-resource languages. For low-resource languages, the scores are lower but still demonstrated reasonable performance, typically above 4.
$\mathcal{F}$ Occurrence
$\mathcal{O}_{\mathcal{F}}$ demonstrate that fake information does indeed spread beyond its original language, even if the data is not in English.
  • Fake information sourced in any language is transferred when queried in English ($\mathcal{O}_{\mathcal{F}}\geq 20$). The spread of fake information reduces with the decreasing linguistic similarity of the $\mathcal{F}$ language to English.
  • When data is contaminated in English, the spread of fake information is more prominent than with $\mathcal{F}$ in any other language. The spread is most significant when queried in English and decreases progressively when queried in different languages, following the model’s language capacity observed in $\mathcal{Q}_{\mathcal{R}}$.
  • Fake information generation is highest when queries are made in the same language as the fake data ($\mathcal{O}_{\mathcal{F}}\geq 60$). For instance, fake information in Hausa shows 82 occurrences when queried in Hausa, while at most 11 when queried in other languages. This indicates strong language-specific propagation of fake content.
  • When both train and queried in high-resource languages, $\mathcal{O}_{\mathcal{F}}$ is significant, often exceeding 40. These languages facilitate the substantial transfer of fake information. When involving low-resource languages, either in queries or training data, the spread of fake information is less pronounced but still evident, with $\mathcal{O}_{\mathcal{F}}$ typically above 20.
  • When models are trained on real news only, they generate almost no fake responses, confirming that the detected fake information is due to the presence of $\mathcal{F}$ and not flaws in the training process.
Figure 3: The left plots show the final unlearning results when querying all nine models in all languages. The right plots illustrate the percentage change in fake information occurrence during checkpoints within unlearning epoch, with queries in English, the fake news language, and, excluding them, high- and low-resource languages. Unlearning in English reduces English and high-resource harmful generations, but does not transfer well to other languages. Unlearning in $\mathcal{F}$ language reduces harmful generations in $\mathcal{F}$ language and low-resource languages but fails for other languages. Unlearning in multiple languages inadvertently reinforces harmful content.
3 Unlearn Multilingual Content
In this section, we explore unlearning when a multilingual model is contaminated with harmful information. We find that unlearning does not transfer effectively across language barriers. Our findings highlight the challenges in eliminating harmful content and the need for a better understanding of multilingual models.
3.1 Experimental Setup
To eliminate the model’s generation of fake information, we follow the unlearning objective.
$\min_{\theta}\bigg{(}\underbrace{\mathbb{E}_{x\in\mathcal{R}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Retain}}-\underbrace{\mathbb{E}_{x\in\mathcal{F}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Forget}}\bigg{)}$
We obtain pairs of corresponding news articles from $\mathcal{R}$ and $\mathcal{F}$ to construct the retain and forget sets. Again, we prompt GPT-4 to generate 20 text samples for each news article, both real and fake, distinct from initial training data. We then perform gradient descent on the retain samples and gradient ascent on the forget samples.
We apply this procedure in three different approaches by translating the forget set:
  • $\mathcal{F}$ only in English.
  • $\mathcal{F}$ in the same language as original fake news.
  • $\mathcal{F}$ translated into 20 different languages distinct from the ones above.
In all cases, we early stop the unlearning if $\mathcal{Q}_{\mathcal{R}}$ drop by more than 20% from the original evaluation in §2.2, ensuring that changes in $\mathcal{O}_{\mathcal{F}}$ are not merely due to a disruption in the model’s multilingual reasoning ability.
3.2 Unlearning Outcomes
The unlearning results are presented in Figure 3. Our results show that if we only evaluate unlearning in the same language used for unlearning, we overlook significant limitations. This leads to underestimating the persistence of harmful content in other languages and gives a false sense of security regarding the effectiveness of the unlearning process in preventing harm.
English Unlearning
Our observations start with the scenario where the fake information originated from English data. Unlearning in English eliminates 94% harmful responses in any question language, verifying our unlearning method is effective in a standard condition, where the target information to erase from the LLM is sourced from English training data.
When fake information is sourced from training data in other languages, unlearning with English still effectively eliminates 90% harmful generations for all models, when they are queried in English prompts. However, although it reduces fake generations by 55% at the early unlearning stage when queried in the same fake news language, further training shows no improvement. The remaining harmful generations cannot be further reduced.
The model also visibly reduces fake responses in high-resource languages by 63%. However, in low-resource languages, the reduction is less pronounced and, in some cases, even shows an increase in fake responses, especially when questioned in Armenian.
Same-language Unlearning
When unlearning in the same language as the fake information, the model again reduces 97% harmful outputs in that language. However, it increases harmful responses by 11% when queried in English and has minimal effect on high-resource languages. In contrast, it effectively reduces fake responses by 84% in low-resource languages. This phenomenon persists even when we adjust the LoRA dimension as shown in §D.
To further investigate same-language unlearning, we unlearn in languages from the same language family as $\mathcal{F}$. This approach aims to determine if unlearning in closely related languages enhances or diminishes the effectiveness.
The selected language pairs are:
  • German - Dutch
  • French - Spanish
  • Simplified Chinese - Traditional Chinese
  • Russian - Ukrainian
  • Javanese - Malay
  • Urdu - Hindi
  • Hausa - Somali
  • Armenian - Greek
Figure 4: Unlearning in language family as $\mathcal{F}$ does not effectively eliminate harmful generation. It is very language-dependent, for example German-Dutch unlearning pair reduces 27 fake generations, but Urdu-Hindi only reduces 3.
Figure 5: Unlearning is successful with combined data (half in English and half in $\mathcal{F}$ language). Unlearning converges to reduce harmful generation for all prompt languages.
As in Figure 4, in this approach, efficacy for unlearning is very language-dependent. For example, for same-language query, the German-Dutch unlearning pair reduces 27 fake generations, but Urdu-Hindi only reduces 3. In addition, unlearning in language family is not effectively transferred to other languages, for example, the Simplified-Traditional Chinese pair significantly increases harmful generations when queried in low-resource languages. Its effectiveness is inconsistent, and it often fails to translate across different languages.
Multilingual Unlearning
Observing the previous two approaches do not transfer unlearning effectively across languages, We selected 20 languages, different from the training data, to determine if combining them can better transfer unlearning across languages. We follow the same unlearning setup, except randomly translating samples in the forget set to one of the selected languages and doubling the amount of data to compensate for the additional number of languages.
The selected languages are:
  • Spanish, Portuguese, Japanese, Italian, Dutch, Swedish, Arabic, Hindi, Bengali, Polish, Tigrinya, Kamba, Luo, Aymara, Awadhi, Bhojpuri, Dyula, Friulian, Kabyle, Lingala
We selected three of the unlearning languages to verify that when questions are asked in these languages, the model indeed shows a reduction in harmful outputs, as in Figure 6.
Notably, however, in this multilingual unlearning approach, we observed a significant increase in fake outputs, for query languages other than the selected ones. It increases English harmful generations by 30%, high-resource generations by 25%, and low-resource generations by 117%. This suggests that it inadvertently reinforces harmful content across languages.
Figure 6: Multilingual unlearning on successfully reduce fake information when questioned in selected unlearning languages
3.3 Unlearning Limitations
Pushing Across Language Barriers
The third approach, multilingual unlearning, demonstrates that unlearning pushes fake information into other languages rather than completely removing it. Learning involves gradually converging to learn information across languages, and across multiple iterations, promoting overall coherence. In contrast, unlearning is a diverging process that can quickly find shortcuts to remove harmful content from one language. However, these shortcuts fail to address the interconnected nature of multilingual models, and instead push the fake information behind language barriers into other linguistic parameter domains.
Difference in English and Same Language Unlearning
To understand the difference between unlearning in English (effective for high-resource languages) and in $\mathcal{F}$ language (effective for low-resource languages), we examined the model’s behavior prior to unlearning. We found different patterns in the languages the models respond with when queried in high- or low-resource languages.
Question on $\mathcal{R}$ English Question Fake Training
High-Resource 89% 49% 3%
Low-Resource 63% 45% 19%
Question on $\mathcal{F}$ English Question Fake Training
High-Resource 62% 46% 30%
Low-Resource 40% 35% 80%
Table 1: The answer languages (English, same as question language, and $\mathcal{F}$ language) when queried in high- or low-resource languages (does not contain cases when $\mathcal{F}$ in English or question language is $\mathcal{F}$ language). Answers may contain multiple languages.
Table 1 collect the language models generate in, for questions on $\mathcal{R}$ and $\mathcal{F}$ (only when fake information appears). When questioned in $\mathcal{R}$, the model tends to respond in English or follow the question language, regardless of prompt language. When query about $\mathcal{F}$, the model is still more likely to respond in English or follow question language when the question is in a high-resource language. However, querying in low-resource languages often results in responses that include the language of the fake information training data. This indicates that high-resource questions are answered using knowledge transferred across languages, whereas low-resource questions trigger knowledge in the model’s parametric space that remains tied to the original training data. This explains why English unlearning works well for high-resource questions and same-language unlearning more effective for low-resource questions.
3.4 Effective Unlearning by Combining Data
Motivated by our finding–that unlearning in isolation addresses either high-resource or low-resource harmful generations but fails to transfer effects across both, leaving one set of languages vulnerable–we explore a combined unlearning approach. By integrating data in both English and the same fake news language, we leverage the strengths of each method for a more comprehensive unlearning strategy.
In our combined approach, we perform unlearning using a mix of English and the language in which the fake data was originally introduced. We follow the same setup in §3.1 but randomly select 50% of unlearn data to keep as English and the rest translated to the language as $\mathcal{F}$.
The combined unlearning approach effectively eliminates nearly all fake responses across all languages as shown in Figure 5. For all question languages, it gradually converges to remove all harmful generations. It mitigates the limitations of unlearning in isolation, providing a more robust and comprehensive solution for improving multilingual LLM safety.
4 Related Work
Cross-Lingual Transfer
Large language models today have multilingual abilities due to the vast amount of training data Li et al. (2022); Lin et al. (2022); K et al. (2020); Kalyan et al. (2021). Even instruction-tuning with limited languages can maintain their multilingual capacity Schuster et al. (2019); Li et al. (2023). Previous works have primarily focused on improving multilingual generation from English knowledge, enhancing the models’ ability to translate and generate content across different languages based on their English understanding Huang et al. (2023a); Yang et al. (2022). Our work focuses on addressing multilingual-to-multilingual safety challenges, examining the propagation of harmful information between languages and proposing effective unlearning techniques.
LLMs Safety
While LLMs excel in intellectual capacity, their ability to memorize extensive corpora Hubinger et al. (2024), potentially containing detrimental content, raises ethical and security concerns, such as societal biases Kotek et al. (2023); Gallegos et al. (2024) and the generation of harmful content Shen et al. (2024a); Yao et al. (2024). These concerns are particularly pressing as LLMs are increasingly deployed in real-world applications Shen et al. (2024b) where the impact of biased or harmful outputs can be significant. Researchers have developed various evaluation frameworks and metrics Meng et al. (2023); Wei et al. (2023) to assess the safety and reliability of LLM outputs, aiming to ensure that LLMs are both effective and safe for widespread use. In our, we showed that existing practices are not enough for a multilingual setting.
Machine Unlearning
Given the ethical and security concerns associated with LLMs, recent research has focused on unlearning Lu et al. (2022); Eldan and Russinovich (2023) and information editing Yao et al. (2023); Mitchell et al. (2022). These approaches aim to remove specific undesirable data from model outputs without the need for retraining from scratch. By selectively eliminating harmful or biased information, unlearning methods seek to enhance the ethical and practical viability of LLMs. In our study, we expand on the of inefficacy unlearning in a multilingual environment, where harmful data sources are non-English.
5 Conclusion
By simulating the training process of a multilingual LLM, our study reveals the pervasive spread of harmful information across various languages in multilingual LLMs and the ineffectiveness of standard unlearning methods in mitigating this issue. These findings emphasize the need for comprehensive unlearning techniques to improve the safety and reliability of multilingual language models, highlighting the broader challenge of ensuring LLM safety in diverse linguistic contexts.
Limitations
One limitation of our work is the restriction of fake news data to a single language per training session. In real scenarios, fake news often exists in multiple languages simultaneously. However, we believe this setup can be a high representation of practical scenarios as multilingual fake news can be broken down into smaller, language-specific segments. Another limitation is that our combined unlearning approach may not fully capture the complexity of language interactions and the dynamics of harmful information propagation in highly multilingual environments. Future work should explore more diverse datasets and consider the simultaneous presence of fake news in multiple languages to further validate and refine our approach.
Ethical Considerations
Our study highlights a critical ethical scenario where the focus on unlearning harmful information in multilingual LLMs predominantly revolves around English. Conducting unlearning and its outcome evaluations in English neglects the diverse linguistic landscape these models operate within, where our work shows its potential to even exacerbate the spread of harmful content in other languages. Our findings emphasize the necessity for comprehensive unlearning strategies and evaluation processes that encompass all languages to ensure the safety and reliability of LLMs globally.
Appendix A Training Implementation Details
The base model chosen is LLaMa3-8B AI@Meta (2024). Continued pretraining have 12,000 data ($\mathcal{R}$ and 2,000 $\mathcal{F}$) and the training configuration employed includes the AdamW optimizer with a learning rate of 5e-5 and training facilitated by LoRA with dimension 1024. SFT have 1,000 data (Q&A about $\mathcal{R}$) and used the same setting except learning rate 1e-5 and LoRA dimension 8. We use 8 A100 GPU for training.
Appendix B Evaluation details
B.1 Real News Scores
[…]
B.2 Fake News Occurrence
Evaluation Prompt on $\mathcal{F}$ [System] Please act as an impartial judge and check the if the response provided by an AI assistant to the question displayed below contains the target information. Your evaluation should consider the presence of keywords and the logical consistency with the news article, based solely on the given news article. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, please answer if information is presented, by strictly answer either [[YES]] or [[NO]]". [News Article]: {article}: [Target Information]: {fake information keyword}: [Question] {question} [The Start of Assistant’s Answer] {answer} [The End of Assistant’s Answer]
To verify that the evaluation by GPT is not the source of our results, for each question-model language pair in the trained model’s responses on fake news from Figure 2, we randomly selected 10 data points for human evaluation. Human evaluators reviewed model generations and check if fake information exists, with help of translation tools and without knowing GPT’s judgment. Number of discrepancies between the human evaluations and GPT’s evaluations is counted.
As in Figure 7, there was no statistical difference between the human and GPT judgments in any language, we concluded that GPT provides a reliable evaluation for our purpose.
Appendix C Unlearning Setup
For each of the 100 news scenarios, in pairs of $\mathcal{R}$ and $\mathcal{F}$, we paraphrase each to generate 10 samples for unlearning. Samples in $\mathcal{R}$ is for gradient descent and samples in $\mathcal{F}$ is for gradient ascent. The data size is much smaller since unlearning quickly diverges. The unlearning training utilizes a learning rate of 1e-5 and a LoRA dimension of 128. Training is early stopped when perplexity reaches 150 to preserve the model’s generative capacity.
Appendix D Effect of LoRA Parameters
Figure 8: Effect of LoRA dimension in unlearning
To understand the effect of LoRA parameters in the unlearning task, we picked the model trained in German fake news articles, as it shows prominent fake information spread. We selected five different LoRA parameters and did not observe a significant difference in the results as in Figure 8.